Busta Limited plans to manufacture bar fridges and the following information is applicable: Estimated sales for the year 5 000 units at R3 400 each Estimated costs for the year: Variable costs Direct Material R520 per unit Direct Labour R350 per unit Variable Manufacturing Cost R110 per unit Selling expenses 6% of selling price per unit sold Factory overheads (all fixed) R625 000 Administrative expenses (all fixed) R462 000 REQUIRED: 1.1 Calculate the total net profit for the estimated figures. (3 marks) 1.2 Calculate the break-even quantity (3 marks) 1.3 Calculate the break-even value (2 marks) 1.4 Calculate the break-even value using the marginal income ratio. (3 marks) 1.5 Calculate the target sales volume to achieve a profit of R920 500.. (3 marks) 1.6 Calculate the new break-even quantity and value if the selling price is increased by 12% (4 marks) 1.7 Calculate the margin of safety in units at the original budgeted volume and price (2 marks)


1
Expert's answer
2021-12-01T12:30:02-0500

Answer


1.1 Net profit = Total Revenue - Total Costs

             Total Revenue = "5000*3400=R17,000,000"

             Total costs = R7,007,000.

           "17,000,000-7,007,000= R9,993,000"

1.2 Break even quantity ="Total fixed costs\over sales per unit-variable costs per unit"

          

Break even quantity ="1,087,000 \over 3400-1184"

 

Break even quantity =490.52 units

 

1.3 Break-even value = "(Break-even quantity) * selling price per unit"


                  "490.52*3400 = R 1,667,768" 490.52*3400 = R1,667,768


1.4 Marginal Income Ratio

     Marginal Income Ratio=     "Price per unit - variable costs \over sales"  

            

Marginal Income Ratio= "2220 \over 3400" = 0.6529


      Marginal Income Ratio= "0.6529 * 100 = 65.29%"


          Marginal Income Ratio = 65.29%.


1.5 Initial profit is 9,993,000 which is when 5000 units are sold. What is the quantity to be sold to achieve 920,000?


"920,000 * 5000 \over 9,993,000"

   = 460 units.

1.6 Initial selling price per unit is 3400 what about when it's increased to 12%.

 "{3400 * 112 \over 100} = 3808"

        

New break even quantity= "{1,087,000 \over 3808 - 1184} = 414.25 Units"


New break even value = "414.25 * 3808 = R1,577,464"


1.7 Margin of Safety = Value of total sales - Break-even value.

       "17,000,000 - 1,667,768 = R15,332,232"